Find the Euclidean distance between (2,5) (4,2)


1
Expert's answer
2021-04-13T07:03:54-0400

Distance "=\sqrt{(Y_2-Y_1)^{2} +(X_2-X_1)^{2} }"


Distance "a" and "b" "=\sqrt{(2-5)^{2} +(4-2)^{2} }"


Distance "a" and "b" "=\sqrt {(-3)^{2} +(2)^{2}}"


Distance ="\sqrt {13}"